# Mutual Information Density of Massive MIMO Systems over Rayleigh-Product Channels

### 05 Jan 2024 v4

@@ C-B1 The evaluation of $U_{1}$
$X_{2}$ $=\mathbb{E}\frac{1}{M}\sum_{i,j}Y^{*}_{j,i}[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}]_{j,i}=\mathbb{E}\{\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}$ (135)
$-\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H}}{M}-\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H}}{M}\}$
=$\overset{(a)}{=}(1-\omega\overline{\omega})\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}-\frac{\omega\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H}}{M}+{\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{\frac{5}{2}}})$
${=}\frac{\overline{\omega}^{2}\operatorname{Tr}{\mathbb{A}}^{2}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M^{2}}+{\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{\frac{5}{2}}}),$
@@
$=\frac{n^{2}\overline{\omega}^{4}\operatorname{Tr}{\mathbb{A}}^{2}}{M^{2}n}\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}+{\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{\frac{5}{2}}})$
=$\overset{(a)}{=}\frac{n^{2}\overline{\omega}^{4}\operatorname{Tr}{\mathbb{A}}^{2}}{M^{2}n}(\frac{1+\delta\overline{\omega}}{1+\delta\overline{\omega}^{2}}\frac{M\omega^{2}}{L}+\frac{M\omega\mathbb{E}\operatorname{Tr}{\mathbb{Q}}^{2}{\mathbb{Z}}{\mathbb{Z}}^{H}}{L\delta(1+\delta\overline{\omega}^{2})M})+{\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{\frac{5}{2}}})$
$=\frac{n^{2}\overline{\omega}^{4}\operatorname{Tr}{\mathbb{A}}^{2}}{M^{2}n}(\frac{1+\delta\overline{\omega}}{1+\delta\overline{\omega}^{2}}\frac{M\omega^{2}}{L}-\frac{M\omega\omega^{\prime}}{L\delta(1+\delta\overline{\omega}^{2})})+{\mathcal{O}}(\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{\frac{5}{2}}}),$
@@
$U_{1,2}=\frac{\rho\operatorname{Tr}{\mathbb{A}}^{2}\overline{\omega}^{4}(\frac{1+\delta\overline{\omega}}{1+\delta\overline{\omega}^{2}}\frac{M\omega^{2}}{L}-\frac{\omega\omega^{\prime}}{\delta(1+\delta\overline{\omega}^{2})})}{M}\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ (137)
$+{\mathcal{O}}(\frac{\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{3}{2}}}+\frac{u\operatorname{Tr}{\mathbb{A}}^{2}}{M^{\frac{3}{2}}}+\frac{u^{2}(1+\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M})}{M}+\frac{u^{3}(1+\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M})}{M^{2}}).$
@@ C-B2 The evaluation of $U_{2}$
$\mathbb{E}\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=\mathbb{E}[\frac{1}{M}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}+\frac{2\sigma^{2}n}{Mn}\operatorname{Tr}{\mathbb{Q}}^{2}{\mathbb{H}}\frac{{\mathbb{X}}{\mathbb{X}}^{H}}{n}{\mathbb{H}}^{H}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$ (138)
=$\overset{(a)}{=}\mathbb{E}[\frac{N}{M}-\frac{\sigma^{2}\operatorname{Tr}{\mathbb{Q}}}{M}-\frac{\sigma^{2}\operatorname{Tr}{\mathbb{Q}}^{2}{\mathbb{H}}{\mathbb{H}}^{H}}{M}+\frac{2\sigma^{2}n}{Mn}\operatorname{Tr}{\mathbb{Q}}^{2}{\mathbb{H}}\frac{{\mathbb{X}}{\mathbb{X}}^{H}}{n}{\mathbb{H}}^{H}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}$
$=\mathbb{E}[\frac{N}{M}-\frac{\sigma^{4}}{M}\operatorname{Tr}{\mathbb{Q}}^{2}-\frac{2\sigma^{2}}{M}\operatorname{Tr}{\mathbb{Q}}^{2}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=(\frac{N}{M}+\frac{\sigma^{4}\delta^{\prime}}{\kappa})\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(\frac{\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{3}{2}}}),$
@@
$\mathbb{E}(\jmath\mu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}-u\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}=(\jmath\sqrt{nM}\overline{C}(\sigma^{2})-uV_{n})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}(u)+\varepsilon({\mathbb{A}},u),$ (139)
@@
$\varepsilon({\mathbb{A}},u)={\mathcal{O}}(f({\mathbb{A}},u,\sigma^{2})+\frac{u}{M}+\frac{(1+u)\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{3}{2}}}+\frac{u\operatorname{Tr}{\mathbb{A}}^{2}}{M^{\frac{3}{2}}}+\frac{u^{2}(1+\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M})}{M}+\frac{u^{3}(1+\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M})}{M^{2}})$ (140)
C-C Step 3: Convergence of $\mathbb{E}e^{\jmath\frac{\gamma_{n}^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}-\overline{\gamma}_{n}}{\sqrt{V_{n}}}}$
By the evaluations in Appendix C-B, we have the following approximation
$\frac{\partial\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}(u)}{\partial u}=\mathbb{E}(\jmath\mu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}-u\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(\frac{u^{2}}{M}+\frac{u^{3}\mathcal{M}(\underline{\nu},u)}{M^{2}}+\frac{u^{4}\mathcal{M}(\underline{\nu},u)}{M^{3}})$ (141)
$=(\jmath\sqrt{nM}\overline{C}(\sigma^{2})-uV_{n})\mathbb{E}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}+{\mathcal{O}}(\frac{u^{2}}{M}+\frac{u^{3}\mathcal{M}(\underline{\nu},u)}{M^{2}}+\frac{u^{4}\mathcal{M}(\underline{\nu},u)}{M^{3}})+\varepsilon({\mathbb{A}},u)$ (142)
$=(\jmath\sqrt{nM}\overline{C}(\sigma^{2})-uV_{n})\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}(u)+{\mathcal{O}}(\frac{u\mathcal{M}(\underline{\nu},u)}{M}+\frac{uV_{n}\mathcal{M}(\underline{\nu},u)}{M^{2}}+\frac{u^{2}}{M}+\frac{u^{3}\mathcal{M}(\underline{\nu},u)}{M^{2}}+\frac{u^{4}\mathcal{M}(\underline{\nu},u)}{M^{3}})+\varepsilon({\mathbb{A}},u)$ (143)
@@
where (141) and (142) follow from Lemma 88 and (139), respectively. (143) is obtained by (95). Solving $\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}(u)$ from the differential equation in (144), we can obtain
$\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}(u)=e^{\jmath u\sqrt{nM}\times\overline{C}(\sigma^{2})-\frac{u^{2}V_{n}}{2}}(1+\int_{0}^{u}e^{-\jmath v\sqrt{nM}\times\overline{C}(\sigma^{2})+\frac{v^{2}V_{n}}{2}}\overline{\varepsilon}({\mathbb{A}},v)\mathrm{d}v).$ (145)
@@
the error term $E(u,{\mathbb{A}})$ can be bounded as
$E(u,{\mathbb{A}})={\mathcal{O}}(\frac{\mathcal{M}(V_{n},u)}{M}+\frac{V_{n}\mathcal{M}(V_{n},u)}{M^{2}}+\frac{u\mathcal{M}(V_{n},u)}{M}+\frac{u^{2}\mathcal{M}(V_{n},u)}{M^{2}}+\frac{u^{3}\mathcal{M}(V_{n},u)}{M^{3}}$ (147)
$+\frac{u}{M}+\frac{u\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{3}{2}}}+\frac{u\mathcal{M}(V_{n},u)(M+\sqrt{\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}})}{M^{2}}+\frac{u^{2}\mathcal{M}(V_{n},u)(M+\sqrt{\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}})}{M^{3}}$
$\frac{\mathcal{M}(V_{n},u)}{M}+\frac{u\sqrt{\operatorname{Tr}{\mathbb{A}}^{2}}}{M^{\frac{3}{2}}}+[\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M^{\frac{3}{2}}}+\frac{u(1+\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M})}{M}+\frac{u^{2}(1+\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M})}{M^{2}}]\mathcal{M}(V_{n},u)).$
Therefore, denoting $\overline{\gamma}_{n}=\sqrt{nM}\overline{C}(\sigma^{2})$, the characteristic function of the normalized version of $\gamma_{n}$ can be written as
$\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}_{norm}(u)=\mathbb{E}e^{\jmath u\frac{\gamma_{n}-\overline{\gamma}_{n}}{\sqrt{V_{n}}}}=\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}(\frac{u}{\sqrt{V_{n}}})e^{-\jmath u\frac{\overline{\gamma}_{n}}{\sqrt{V_{n}}}}+E(\frac{u}{\sqrt{V_{n}}},{\mathbb{A}})\overset{(a)}{=}e^{-\frac{u^{2}}{2}}+{\mathcal{O}}(\frac{1}{\sqrt{M}}),$ (148)
where step $(a)$ in (148) follows from $\mathcal{M}(V_{n},u)=\frac{\mathcal{M}(1,\frac{u}{\sqrt{V_{n}}})}{V_{n}}$ and
${\mathcal{O}}(E(\frac{u}{\sqrt{V_{n}}},,{\mathbb{A}}))={\mathcal{O}}(\frac{\mathcal{M}(u)\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}}{\sqrt{M}V_{n}})={\mathcal{O}}(\frac{\mathcal{M}(u)\frac{\operatorname{Tr}{\mathbb{A}}^{2}}{M}}{\sqrt{M}(\frac{\tau\operatorname{Tr}({\mathbb{A}}^{2})}{M}+K)})={\mathcal{O}}(\frac{\mathcal{M}(1,u)}{\sqrt{M}}),$ (149)
and $\tau={\mathcal{O}}(1)$ is the coefficient of $\frac{\operatorname{Tr}{\mathbb{A}}_{n}^{2}}{M}$ in (16) and $K$ does not depend on $M$, $N$, $L$, and $n$. By (148) and Lévy’s continuity theorem, there holds true that

### 09 Apr 2024 v5

@@ C-B1 The evaluation of $U_{1}$
$X_{2}$ $=\frac{1}{M}\sum_{i,j}\mathbb{E}[Y^{*}_{j,i}[{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}]_{j,i}]=\mathbb{E}\Bigl{[}\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})}{M}$ (135)
$-\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})}{M}\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H})}{M}-\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})}{M}\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H})}{M}\Bigr{]}$
=$\overset{(a)}{=}(1-\omega\overline{\omega})\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}-\frac{\omega\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{A}}^{2}{\mathbb{H}}^{H})]}{M}+{\mathcal{O}}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{\frac{5}{2}}}\Bigr{)}$
${=}\frac{\overline{\omega}^{2}\operatorname{Tr}({\mathbb{A}}^{2})\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})]}{M^{2}}+{\mathcal{O}}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{\frac{5}{2}}}\Bigr{)},$
@@
$=\frac{n^{2}\overline{\omega}^{4}\operatorname{Tr}({\mathbb{A}}^{2})}{M^{2}n}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}+{\mathcal{O}}(\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{\frac{5}{2}}})$
=$\overset{(a)}{=}\frac{n^{2}\overline{\omega}^{4}\operatorname{Tr}({\mathbb{A}}^{2})}{M^{2}n}\Bigl{(}\frac{1+\delta\overline{\omega}}{1+\delta\overline{\omega}^{2}}\frac{M\omega^{2}}{L}+\frac{M\omega\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}^{2}{\mathbb{Z}}{\mathbb{Z}}^{H})]}{L\delta(1+\delta\overline{\omega}^{2})M}\Bigr{)}+{\mathcal{O}}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{\frac{5}{2}}}\Bigr{)}$
$=\frac{n^{2}\overline{\omega}^{4}\operatorname{Tr}({\mathbb{A}}^{2})}{M^{2}n}\Bigl{(}\frac{1+\delta\overline{\omega}}{1+\delta\overline{\omega}^{2}}\frac{M\omega^{2}}{L}-\frac{M\omega\omega^{\prime}}{L\delta(1+\delta\overline{\omega}^{2})}\Bigr{)}+{\mathcal{O}}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{\frac{5}{2}}}\Bigr{)},$
@@
$U_{1,2}=\frac{\rho\operatorname{Tr}({\mathbb{A}}^{2})\overline{\omega}^{4}(\frac{1+\delta\overline{\omega}}{1+\delta\overline{\omega}^{2}}\frac{M\omega^{2}}{L}-\frac{\omega\omega^{\prime}}{\delta(1+\delta\overline{\omega}^{2})})}{M}\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]$ (137)
$+{\mathcal{O}}\Bigl{(}\frac{\sqrt{\operatorname{Tr}({\mathbb{A}}^{2}})}{M^{\frac{3}{2}}}+\frac{u\operatorname{Tr}({\mathbb{A}}^{2})}{M^{\frac{3}{2}}}+\frac{u^{2}(1+\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M})}{M}+\frac{u^{3}(1+\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M})}{M^{2}}\Bigr{)}.$
@@ C-B2 The evaluation of $U_{2}$
$\mathbb{E}[\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=\mathbb{E}\Bigl{[}[\frac{1}{M}\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}{\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})+\frac{2\sigma^{2}n}{Mn}\operatorname{Tr}\Bigl{(}{\mathbb{Q}}^{2}{\mathbb{H}}\frac{{\mathbb{X}}{\mathbb{X}}^{H}}{n}{\mathbb{H}}^{H}\Bigr{)}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}\Bigr{]}$ (138)
=$\overset{(a)}{=}\mathbb{E}\Bigl{[}[\frac{N}{M}-\frac{\sigma^{2}\operatorname{Tr}({\mathbb{Q}})}{M}-\frac{\sigma^{2}\operatorname{Tr}({\mathbb{Q}}^{2}{\mathbb{H}}{\mathbb{H}}^{H})}{M}+\frac{2\sigma^{2}n}{Mn}\operatorname{Tr}\Bigl{(}{\mathbb{Q}}^{2}{\mathbb{H}}\frac{{\mathbb{X}}{\mathbb{X}}^{H}}{n}{\mathbb{H}}^{H}\Bigr{)}]\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}\Bigr{]}$
$=\mathbb{E}\Bigl{[}\frac{N}{M}-\frac{\sigma^{4}}{M}\operatorname{Tr}({\mathbb{Q}}^{2})-\frac{2\sigma^{2}}{M}\operatorname{Tr}({\mathbb{Q}}^{2}{\mathbb{H}}{\mathbb{A}}{\mathbb{H}}^{H})\Bigr{]}\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=(\frac{N}{M}+\frac{\sigma^{4}\delta^{\prime}}{\kappa})\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+{\mathcal{O}}\Bigl{(}\frac{\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{M^{\frac{3}{2}}}\Bigr{)},$
@@
$\mathbb{E}[(\jmath\mu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}-u\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]=(\jmath\sqrt{nM}\overline{C}(\sigma^{2})-uV_{n})\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+\varepsilon({\mathbb{A}},u),$ (139)
@@
$\varepsilon({\mathbb{A}},u)={\mathcal{O}}\Bigl{(}f({\mathbb{A}},u)+\frac{u}{M}+\frac{(1+u)\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{M^{\frac{3}{2}}}+\frac{u\operatorname{Tr}({\mathbb{A}}^{2})}{M^{\frac{3}{2}}}+\frac{u^{2}(1+\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M})}{M}+\frac{u^{3}(1+\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M})}{M^{2}}\Bigr{)}$ (140)
C-C Step 3: Convergence of $\mathbb{E}\Bigl{[}e^{\jmath\frac{\gamma_{n}^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}-\overline{\gamma}_{n}}{\sqrt{V_{n}}}}\Bigr{]}$
By the evaluations in Appendix C-B, we can obtain the evaluation for in (141)-(144) given by
$\frac{\partial\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}(u)}{\partial u}=\mathbb{E}[(\jmath\mu^{{\mathbb{Y}},{\mathbb{Z}}}_{n}-u\nu^{{\mathbb{Y}},{\mathbb{Z}}}_{n})\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+{\mathcal{O}}\Bigl{(}\frac{u^{2}}{M}+\frac{u^{3}\mathcal{M}(u)}{M^{2}}+\frac{u^{4}\mathcal{M}(u)}{M^{3}}\Bigr{)}$ (141)
$=(\jmath\sqrt{nM}\overline{C}(\sigma^{2})-uV_{n})\mathbb{E}[\Phi_{n}^{{\mathbb{Y}},{\mathbb{Z}}}]+{\mathcal{O}}\Bigl{(}\frac{u^{2}}{M}+\frac{u^{3}\mathcal{M}(u)}{M^{2}}+\frac{u^{4}\mathcal{M}(u)}{M^{3}}\Bigr{)}+\varepsilon({\mathbb{A}},u)$ (142)
$=(\jmath\sqrt{nM}\overline{C}(\sigma^{2})-uV_{n})\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}(u)+{\mathcal{O}}\Bigl{(}\frac{u\mathcal{M}(u)}{M}+\frac{uV_{n}\mathcal{M}(u)}{M^{2}}+\frac{u^{2}}{M}+\frac{u^{3}\mathcal{M}(u)}{M^{2}}+\frac{u^{4}\mathcal{M}(u)}{M^{3}}\Bigr{)}+\varepsilon({\mathbb{A}},u)$ (143)
@@
where (141) and (142) follow from (88) and (139), respectively. (143) is obtained by (95). Solving $\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}(u)$ from the differential equation in (144), we can obtain
$\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}(u)$ $=e^{\jmath u\sqrt{nM}\times\overline{C}(\sigma^{2})-\frac{u^{2}V_{n}}{2}}(1+\int_{0}^{u}e^{-\jmath v\sqrt{nM}\times\overline{C}(\sigma^{2})+\frac{v^{2}V_{n}}{2}}\overline{\varepsilon}({\mathbb{A}},v)\mathrm{d}v)$ (145)
$=e^{\jmath u\sqrt{nM}\times\overline{C}(\sigma^{2})-\frac{u^{2}V_{n}}{2}}+E(u,{\mathbb{A}}).$
@@
the error term $E(u,{\mathbb{A}})$ can be bounded by
$E(u,{\mathbb{A}})={\mathcal{O}}\Bigl{(}\frac{\mathcal{M}(V_{n},u)}{M}+\frac{V_{n}\mathcal{M}(V_{n},u)}{M^{2}}+\frac{u\mathcal{M}(V_{n},u)}{M}+\frac{u^{2}\mathcal{M}(V_{n},u)}{M^{2}}+\frac{u^{3}\mathcal{M}(V_{n},u)}{M^{3}}$ (147)
$+\frac{u}{M}+\frac{u\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{M^{\frac{3}{2}}}+\frac{u\mathcal{M}(V_{n},u)\Bigl{(}M+\sqrt{\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}}\Bigr{)}}{M^{2}}+\frac{u^{2}\mathcal{M}(V_{n},u)\Bigl{(}M+\sqrt{\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}}\Bigr{)}}{M^{3}}$
$+\frac{\mathcal{M}(V_{n},u)}{M}+\frac{u\sqrt{\operatorname{Tr}({\mathbb{A}}^{2})}}{M^{\frac{3}{2}}}+\Bigl{[}\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M^{\frac{3}{2}}}+\frac{u(1+\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M})}{M}+\frac{u^{2}(1+\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M})}{M^{2}}\Bigr{]}\mathcal{M}(V_{n},u)\Bigr{)}.$
Denoting $\overline{\gamma}_{n}=\sqrt{nM}\overline{C}(\sigma^{2})$, the characteristic function of the normalized version of $\gamma_{n}$ can be written as
$\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}_{norm}(u)$ $=\mathbb{E}e^{\jmath u\frac{\gamma_{n}-\overline{\gamma}_{n}}{\sqrt{V_{n}}}}=\Psi^{{\mathbb{W}},{\mathbb{Y}},{\mathbb{Z}}}\Bigl{(}\frac{u}{\sqrt{V_{n}}}\Bigr{)}e^{-\jmath u\frac{\overline{\gamma}_{n}}{\sqrt{V_{n}}}}+E\Bigl{(}\frac{u}{\sqrt{V_{n}}},{\mathbb{A}}\Bigr{)}\overset{(a)}{=}e^{-\frac{u^{2}}{2}}+{\mathcal{O}}\Bigl{(}\frac{1}{\sqrt{M}}\Bigr{)},$ (148)
where step $(a)$ in (148) follows from
${\mathcal{O}}\Bigl{(}E(\frac{u}{\sqrt{V_{n}}},{\mathbb{A}})\Bigr{)}={\mathcal{O}}\Biggl{(}\frac{\mathcal{M}(u)\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}}{\sqrt{M}V_{n}}\Biggr{)}={\mathcal{O}}\Biggl{(}\frac{\mathcal{M}(u)\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}}{\sqrt{M}(\frac{\tau\operatorname{Tr}({\mathbb{A}}^{2})}{M}+K)}\Biggl{)}={\mathcal{O}}\Bigl{(}\frac{\mathcal{M}(u)}{\sqrt{M}}\Bigr{)},$ (149)
and $\tau={\mathcal{O}}(1)$ is the coefficient of $\frac{\operatorname{Tr}({\mathbb{A}}^{2})}{M}$ in (16) and $K$ does not depend on $M$, $N$, $L$, and $n$. By (148) and Lévy’s continuity theorem [52], there holds true that